OpenStudy (anonymous):

Exponential functions: y=-2^x+3 finding the domain, range, and intercepts

OpenStudy (anonymous):

domain is all real number because you can raise 2 to any power you like

OpenStudy (anonymous):

\[f(x)=-2^x+3\] range will be \[(-\infty,0)\]

OpenStudy (anonymous):

but \[2^x\geq 0 \] for all x so \[-2^x\leq 0\] and therefore \[-2^x+3\leq 3\] for any x

OpenStudy (anonymous):

oooops

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

I am rusty

OpenStudy (anonymous):

agree; totally.

OpenStudy (anonymous):

now for intercepts. y intercept put \[x=0\] get \[-2^0+3=-1+3=2\] so (0,2) is y intercept

OpenStudy (anonymous):

Range is all y values strictly less than 3. There is an horizontal symptom at y=3.

OpenStudy (anonymous):

for x intercept set \[-2^x+3=0\] and solve you get \[2^x=3\] so \[x=\log_3(2)=\frac{\ln(3)}{\ln(2)}\]

OpenStudy (anonymous):

last one in case you need a decimal you can use that and a calculator to find it

OpenStudy (anonymous):

how would you solve sin^2 (pi/4) - cos^2 (pi/3)

OpenStudy (anonymous):

Thank you so much I was basically there but needed to finish it up.

OpenStudy (anonymous):

....................

OpenStudy (anonymous):

can someone help me PLEASE?